OpenStudy (anonymous):

Problem involving magnetism and current! Any help would be greatly appreciated! :)

OpenStudy (anonymous):

Many living things can sense magnetic fields and have an innate tendency to face magnetic north regardless of wind strength, time of day, or position of the sun. In an experiment, a gold ring ID collar of radius 0.06m was attached on a fox. It takes the fox 0.2s to move from a region in which the collar is perpendicular to earth's magnetic field of 5.0E-5 T to one in which the collar is parallel to Earth's magnetic field.. What current, if any, is induced in the collar?

OpenStudy (anonymous):

@Michele_Laino @IrishBoy123

OpenStudy (vincent-lyon.fr):

Use Faraday's law of induction.

OpenStudy (anonymous):

@Vincent-Lyon.Fr could you elaborate on that please? I'm not saying give me the answer, but I am completely stuck on how to approach this problem =/

OpenStudy (michele_laino):

the change of cocatenated magnetic flux with the loop is: \[\Delta \Phi = \pi {r^2}B\] where \(r\) is the radius of the loop, and \(B\) is the magnetic field. The time in which such magentic flux change occurred is \( \Delta t=0.2\) seconds. So the electrical current \(I\) of the loop, has the subsequent magnitude: \[I = \frac{1}{R}\frac{{\Delta \Phi }}{{\Delta t}} = \frac{{\pi {r^2}B}}{{R\Delta t}}\] where \(R\) is the electrical resistance of the loop or collar around the fox

OpenStudy (michele_laino):

In order to find \(R\) we can apply this formula: \[R=\rho \frac{{2\pi r}}{s}\] where \(\rho\) is the specific resistance of gold, which is tabulated, and \(s\) is the croos sectional area of the collar

OpenStudy (michele_laino):

cross* sectional area

OpenStudy (michele_laino):

nevertheless, from your data, we aren't to find the value of \(R\) since, we don't know the value of \(s\)

OpenStudy (michele_laino):

oops..we aren't able to find...

OpenStudy (michele_laino):



OpenStudy (anonymous):

Hmm, are you sure that there isn't another way to calculate the current? It seems as if the question is pushing us to find a numerical answer rather than a function. But perhaps not?

OpenStudy (michele_laino):

if we apply the Faraday's law, we are able to find the voltage along the collar, nevertheless, if we want to compute the magnitude of the corresponding current, we need to know the electrical resistance of the collar

OpenStudy (anonymous):

Hmm, what if we were able to consider it as a very thin wire?

OpenStudy (michele_laino):

I think that a collar has to have a cross sectional area

OpenStudy (michele_laino):

oops.. another typo: concatenated*

OpenStudy (anonymous):

Hmm, okay. Maybe it's best to go along the route with Faraday's law and voltage/emf? I know that \[emf=- \frac{ d\Phi_M }{ dt }\]and \[\Phi_M=\int\limits B*dA\]

OpenStudy (michele_laino):

what is the numerical result, please?

OpenStudy (anonymous):

I don't have the answer =/

OpenStudy (michele_laino):

yes! your procedure above is correct, nevertheless, we can apply the Faraday's law in integral form

OpenStudy (anonymous):

How do you mean integral form?

OpenStudy (michele_laino):

I mean without using differentials

OpenStudy (michele_laino):

like this: \[E = \frac{{\Delta \Phi }}{{\Delta t}}\]

OpenStudy (anonymous):

Ah okay. So once I have the emf, then \[V_{emf}=IR\] right? But we're not given R.. hmm

OpenStudy (michele_laino):

yes! Your point of view is right!

OpenStudy (anonymous):

Hmm, if I say \[emf=vBd\]then I could relate the two as\[vBd=-\frac{ B (\Delta A) }{ \Delta t }\]Right? Looks like the B cancels out anyway. But then I don't know v so it looks like I'm still stuck anyway..

OpenStudy (michele_laino):

please the left side is the e.m.f. at the ends of a conductor which is moving inside a magnetic field. Whereas your problem can be modeled by this drawing: 

OpenStudy (anonymous):

Hmm, okay, where can I go after this?

OpenStudy (anonymous):

I see that \(\Phi_M\) would be 0 after, since A is parallel to B, correct?

OpenStudy (michele_laino):

yes! Before the flux is \(\Phi = \pi {r^2}B\), and after the flux is \(\Phi =0\), so its change is: \[\Delta \Phi = 0 - \pi {r^2}B = - \pi {r^2}B\]

OpenStudy (michele_laino):

of course, with the word "flux", I mean the concatenated flux with the loop

OpenStudy (anonymous):

Right, which means that: \[emf=\frac{ \pi r^2B }{ \Delta t }\]

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Okay, so once I have the emf, how can I relate this back to current?

OpenStudy (michele_laino):

we can apply the Ohms's law

OpenStudy (anonymous):

But we don't know R? >_<

OpenStudy (anonymous):

Oh, we would just look that up right?

OpenStudy (michele_laino):

Yes, I know! Nevertheless, the collar is made by gold, which is a metal, more precisely a noble metal, and the conduction inside metals is modeled by the Ohm's law

OpenStudy (anonymous):

I found that \(\rho_{Au}=22.1 \ n\Omega \ m\) So then we can just plug everything in at this point, right?\[\frac{ \pi r^2 B }{ (22.1\times 10^{-9})t }=I\]

OpenStudy (anonymous):

@Michele_Laino I got 127.94 amps -- does this seem reasonable?

OpenStudy (michele_laino):

I think that it is too high

OpenStudy (anonymous):

I thought so too... But everything seems correct. Field is in T, converted from nano-ohms to Ohms, time in s, radius in m...

OpenStudy (michele_laino):

we have to divide by \(R\) not by \(\rho\)

OpenStudy (michele_laino):

please wait a I moment, I want to do a little research on the terrestrial magnetic field

OpenStudy (anonymous):

Well, I suppose I should leave it in terms of S (cross sectional area of the collar) then. Normally our professor would specify if we need to leave it in terms of a variable but I don't see another way to do it.

OpenStudy (michele_laino):

In that case the formula for the requested current, is: \[I = \frac{{rB}}{{2\rho \Delta t}}s\]

OpenStudy (michele_laino):

\[\Huge I = \frac{{rB}}{{2\rho \Delta t}}s\]

OpenStudy (irishboy123):

yes, some info seems to be missing here a fairly simple approach is to take Faraday's law \(\large \mathcal{E} = -{{d\Phi} \over dt} \ \) noting that \(|\Phi| = B . A\) and \(A = \pi r^2 \cos \theta\) ...... where \(\theta\) is the angle between (A) the vector that is normal to the face of the ring and (B) the field ...... which angle varies from \( 0 \; \text{ to } \; \pi/2\) over the 0.2s. so \(\large \mathcal{E} = -{{d\Phi} \over dt} \ = -B \pi R^2 {{d[ \cos \theta]} \over dt} = B \pi R^2\ \sin \theta \, \dot \theta \) assuming \(\dot \theta = const. = \dfrac{\pi/2}{0.2} = \dfrac{5 \pi}{2}\) and inserting resistance as \(\dfrac{\rho 2 \pi R}{a}\) where a is the cross sectional area of the gold ring, seems that to convert this to a current using the known resistivity or conductivity of gold, you do need to know the radius of the ring. \(\large I(\theta) = \dfrac{B \pi R^2\ \sin \theta \, \frac{5 \pi }{2} . a}{\rho 2 \pi R} = \dfrac{5\pi R B }{4 \rho}.a.\sin \theta\)

OpenStudy (anonymous):

@IrishBoy123 I don't understand the significance and inclusion of \(\dot \theta = const. = \dfrac{\pi/2}{0.2} = \dfrac{5 \pi}{2}\) =/

OpenStudy (irishboy123):

it goes from perp to parallel in 0.2 sec. so \(\theta\) goes from 0 to π/2 in that time

OpenStudy (anonymous):

@Michele_Laino Ok so I get about 339.67*S amps. If we randomly choose that S has a radius of 1cm=0.01m, then we have an area of 3.14E-4. (3.14E-4)(339.67)=0.107amps Is this reasonable now? :P

OpenStudy (michele_laino):

Yes, I think so!